OpenStudy (anonymous):

i want to find the inverse laplace transform of the function H(s)=s+7/(s^2-3*s-10) using partial fractions so H(s)= A/(s+2) + B/(s-5). what is A=? what is B=? i want to find the inverse laplace transform of the function H(s)=s+7/(s^2-3*s-10) using partial fractions so H(s)= A/(s+2) + B/(s-5). what is A=? what is B=? @Mathematics

OpenStudy (amistre64):

s+7/(s^2-3*s-10) = A/(s+2) + B/(s-5) ; multiply both side by the common denominator s+7= A(s-5) + B(s+2) now the easiest thing to do is to zero out A then B, say s=5 5+7= A(5-5) + B(5+2) 12= B(7) B = 7/12 say s=-2 -2+7= A(-2-5) + B(-2+2) 5= A(-7) A = -5/7

OpenStudy (anonymous):

b=12/7. now apply the inverse laplace transform to the above result?

OpenStudy (amistre64):

if that is your goal then yes. But I dont know how to do a laplace inverse. its on my todo list :)